OpenStudy (anonymous):

Solve: 3(x + 4) = 5(x − 2)

OpenStudy (anonymous):

x=11

OpenStudy (jackellyn):

Distribute 3(x) + 3(4) = 5(x) + 5(-2) 3x + 12 = 5x -10 12+10 = 5x-3x 22 = 2x x=11

OpenStudy (anonymous):

3(x+4)=5(x-2) 3x+12=5x-10 12+10=5x-3x 22=2x x=11

OpenStudy (anonymous):

3(x + 4) = 5(x − 2) multiply the numbers in the brackets with the outside number on both sides 3x+12=5x-10 5x-3x=12+10 2x=22 x=11 so x=11 is the answer :)